OpenStudy (anonymous):

Which expression is the simplified form of the square root of h to the eleventh power.? h3the square root of h squared. h10the square root of h. h5the square root of h. hthe square root of h to the fifth power.

OpenStudy (campbell_st):

well start by using some index laws \[\sqrt{h} = h^{\frac{1}{2}}\] so \[(\sqrt{h})^{11} = (h^{\frac{1}{2}})^{11}\] multiply the powers \[h^{\frac{11}{2}}\] now writing it in radical form its \[\sqrt{h^{11}}= \sqrt{h^{10} \times h} = \sqrt{h^{10}} \times \sqrt{h} = h^{5} \times \sqrt{h}\] hope it helps

OpenStudy (anonymous):

it did thank you :)